# The IACOB project

### v1

@@ 4.2 Spin-rates of the sample
Another observational characteristic proposed to delineate the location of the TAMS is the drop of the upper envelope of the distribution as a function of (see Vink et al. 2010; Brott et al. 2011). Figure 4 shows our distribution of values against separating stars in the same luminosity ranges as in Fig. 3. To qualitatively evaluate the aforementioned feature, we included all the stars regardless of their distance. The figure indicates the location of the ZAMS (gray areas), the proposed empirical TAMS (dashed colored lines), and outlines the critical rotational velocity (; solid green lines), that is, the equatorial velocity at which the centrifugal force balances gravity, as a function of (see below). We also include the median values for stars within bins of 4 kK (black-edge colored dots) and the associated 25% and 75% (solid black lines) and 10% and 90% (dashed gray lines) percentiles. Although a detailed discussion of the connection with the TAMS is presented in Sect. 5.1.1, here we highlight the three most interesting features of Fig. 4 (see also Appendix C).
A first feature is the separation of stars into two main components (see also Fig. 10 of de Burgos et al. (2024b)): a low- densely populated group of stars (colored differently) and a less populated group of fast-rotating objects (cyan circles; indicated also in Fig. 2). Following de Burgos et al. (2024b), we adopt the value of =100 km s-1 that best separates both components. This separation is well known to be present in the O-star domain and extends toward the early-B supergiant domain (see, e.g., Holgado et al. 2022; de Burgos et al. 2023, and references therein). Using as reference those stars within 2500 pc, we find a decreasing fraction of fast-rotating objects from 50% for stars with luminosities below 4.65 dex, to 35% for stars between 5.0 and 4.65 dex, and to 30% for stars with luminosities higher than 5.0 dex. On average, 38% of the stars located on the hot side of the TAMS belong to the fast-rotating component. A further discussion of both components can be found in Sect. 5.1.2 and Sect. 5.3.
@@
A third characteristic is the consistent decrease of the maximum reached by stars with decreasing within each luminosity range. In fact, the upper envelopes closely follow , which is shown with a solid green line (shown also in the right panel of Fig. 2), and has been calculated from the same models used in Fig. 2 employing Eq. 2 in Langer (1998). All these features are discussed in more detail in Sect. 5.1.
@@ 4.4 Distribution of SB1 objects in the HR diagram
Additionally, observational biases due to the different intrinsic brightnesses of the stars are likely present in their work, whereas our volume-limited sample with high degree of completeness solves this issue. Furthermore, while the sHR diagram is used in Castro et al. (2014), here we make use of the HR diagram, combined with additional information on the spin-rate properties and binary fraction of the stars (Sects. 4.2, and 4.4). Moreover, we consider double the number of stars in Castro et al. (2014) up to 2500 pc within the same range. We also note that in Castro et al. (2014), the surface gravities were not corrected by centrifugal forces (see Herrero et al. 1992; Repolust et al. 2004), shifting fast-rotating stars towards apparently higher luminosities in the sHR diagram.
Despite the differences, Castro et al. (2014) found a drop in the density of stars below 25 M⊙ at 20 kK, which was also interpreted as the TAMS. However, for stars with 25 – 40 M⊙ the drop was located at 10 kK. As shown in Figs. 2 and 3, our results indicate that the drop occurs throughout the full range of luminosities and masses considered here, and is located approximately at the same (24 – 22 kK). The overdensity of stars shown in Castro et al. (2014) beyond 20 kK, which is not observed in any of the panels of Fig. 2, is likely the result of a selection bias introduced by considering any star with available parameters from the literature rather than a volume-limited sample. In particular, since mid-B-type stars are intrinsically brighter in the optical compared to early-B- and O-type objects, the aforementioned overdensity is likely caused by mid-B-type stars located at larger distances.
@@ 5.1 Spin-rate properties of the sample
As indicated in Sect. 4.2, two main observable features are known to characterize the spin rates of luminous O- and B-type stars. One is the bimodal distribution of the projected rotational velocity (e.g. Conti & Ebbets 1977; Howarth et al. 1997; Dufton et al. 2013; Ramírez-Agudelo et al. 2013; Holgado et al. 2022; de Burgos et al. 2024b), which includes a main component of slowly-rotating stars and a secondary more extended tail of fast-rotating objects. The latter have been proposed to be the products of binary interactions (e.g., Ramírez-Agudelo et al. 2013; de Mink et al. 2014; Ramírez-Agudelo et al. 2015; Holgado et al. 2022; Britavskiy et al. 2023; de Burgos et al. 2024b). The second feature is the lack of fast-rotating stars below a certain temperature (see Howarth et al. 1997; Ryans et al. 2002; Vink et al. 2010; Simón-Díaz & Herrero 2014; McEvoy et al. 2015; de Burgos et al. 2024b). As indicated by different studies, this characteristic could be interpreted as the empirical location of the TAMS (see Vink et al. 2010; Brott et al. 2011; Martinet et al. 2021; de Burgos et al. 2024b). We explore this possibility further.
@@ 5.1.1 The lack of fast-rotating stars beyond the TAMS
On the other hand, stars with expanding envelopes and high spin rates near the TAMS can get too close to their critical rotational velocities (see, e.g., Meynet & Maeder 2006; Meynet et al. 2007). It is predicted that before reaching that point, stars expel enough material into the medium (e.g. Kurfürst et al. 2014), triggering a spin-down of the star’s rotation (e.g. Georgy et al. 2011; de Mink et al. 2013). As shown in Fig. 4 (see also the right panel of Fig. 2), the critical rotation that we extract from the models approximates the upper envelope of the distribution, particularly in the two central panels. We note that, except for the top panel, we could, in principle, expect to find some fast-rotating objects that have not yet reached their .
@@
A third possibility is related to mass loss. In this respect, Vink et al. (2010) suggested that the bi-stability nature of the winds (see Pauldrach & Puls 1990; Vink et al. 1999, 2000) can lead to an important increase in the mass-loss rates at 22 kK, and to a subsequent enhancement of the angular momentum loss (known as the “bi-stability braking”). However, in de Burgos et al. (2024a) (see also Rubio-Díez et al. 2022; Krtička et al. 2024; Bernini-Peron et al. 2024) it is shown that such an increase in mass-loss rates (by a factor of ten to twenty) is not observed444Newer prescriptions, such as those of Björklund et al. (2021), do not predict any enhancement near our empirical TAMS.. Therefore, we argue that the lack of fast-rotating stars is not due to the latter proposed mechanism.
@@
Given these considerations, the lack of fast-rotating objects could be the result of the statistically and relatively lower number of stars expected beyond the TAMS, that is, the chances of observing these objects is lower. This could be further supported by the following. Considering that for stars with = 5.35 – 4.65 (where we account for more stars), the histograms of Fig. 3 show the largest difference in the relative number of stars when crossing the TAMS, which we locate at an average of 22.5 kK. For stars up to 2500 pc, there are five times fewer stars between those located on the hot and the cool sides of the TAMS within 5 kK. Considering that the proportion of fast-rotating objects located within each side with respect to the total is 1:3, this exercise would result in only six fast-rotating objects below 22.5 kK. We account for two objects that meet these conditions, which we consider sufficient to support the aforementioned possibility. We note that we do not account for the fact that some of the fast-rotating stars may not be seen as such because of a projection effect. Furthermore, if fast-rotating stars are the product of binary interaction, it is expected to occur mainly when at least one of the components of the binary system is still on the MS (see de Mink et al. 2013, 2014). Therefore, it would be difficult to observe these objects beyond the TAMS. Additionally, fast-rotating objects evolving near or beyond the TAMS will eventually reach their , thus becoming slowly-rotating stars. Nonetheless, one cannot rule out the possibility of unconsidered effects or forces that can trigger moderate braking toward the end of the MS. This can include, for example, the effect of magnetic fields (see Keszthelyi et al. 2020, 2022).
@@ 5.1.2 The evolution of along the MS
Interestingly, the median values of the two top panels in Fig. 4 show similar trends towards lower effective temperatures. The values are approximately constant until 30 kK and then slowly decrease toward the cool end. This trend is less obvious in the third panel, but may still be present, and it is certainly not the case of the bottom panel, which we discuss separately in Appendix E. We also note that the first median value may not be so representative, as it overlaps with the position of the ZAMS, where the number of stars is significantly lower than for the other mean values. Excluding that one, our results strongly suggest a slow-braking process occurring along the MS, which has already been empirically supported by the works of Holgado et al. (2022); de Burgos et al. (2024b). Furthermore, the number of stars on the hot side of the TAMS with 100 km s-1 (i.e., the low- component of the spin-rate distribution) represents 65% of the total number. This large fraction may indicate that massive stars are born with low or mild initial rotations (not more than 0.2 of the critical velocity), as indicated also in the aforementioned works. However, we note that since we do not account for projection effects, that percentage may represent a lower limit.
@@ 5.2 The impact of binaries
In Sect. 4.4 we present the distribution of detected SB1 systems in an HR diagram, limited to a distance of 2500 pc. Figure 6 shows the proposed TAMS and the lack of SB1 systems beyond it, on the cool side. By evaluating the fraction of SB1 systems on each side of the TAMS, we obtain 39% within = 30 – 22.5 kK, and 15% within = 22.5 – 15 kK, highlighting an important decrease when crossing the TAMS. Interestingly, the former percentage is similar to the 367% found by Dunstall et al. (2015) in the 30 Doradus region of the Large Magellanic Cloud. Furthermore, our results approach those found by McEvoy et al. (2015) in the Tarantula Nebula, where SB1 systems were not detected below a given temperature. In this work, we show that SB1 systems are also present beyond the TAMS, but their number is very low. We conclude that the lower percentage of SB1 systems beyond the TAMS can be used to infer its location with respect to . We also argue that the few SB1 systems beyond the TAMS likely correspond to stars heading to the RSg phase, since this phase practically eliminates all detectable binaries beyond the TAMS (see de Mink et al. 2014).
@@
Our results also indicate that the percentage of SB1 systems within = 31 – 22.5 kK decreases in the upper half of the HR diagram. In particular, from 40% for stars with ¡ 5.0 to 31% for stars above that luminosity. This case could be partially explained by the higher detection threshold of RV (up to 25 km s-1; see Appendix D) towards our highest luminosities, which would eliminate some true SB1 systems.
Similarly as for SB1 systems, the presence of a drop in SB2+ systems on the cool side of the TAMS could provide additional evidence of its location. Although SB2+ systems were intentionally excluded from our sample, and therefore we do not count with reliable quantitative spectroscopic analysis of each individual component, we can still use to some extent the available information on their spectral types as a proxy of the corresponding effective temperatures.
@@
These results also provide evidence that the number of SB2+ systems in the BSG domain before reaching the TAMS is lower than for O9 stars. In this regard, given their relatively low number, the lack of SB2+ systems in the sample should not produce important biases with respect to the distribution of stars in Figs. 2, 3. However, the difference between O9- and B0-type stars could partially explain the apparent lower density of objects around 32 kK (see Sect. 4.1). We also notice the larger number of B3 – B4 SB2+ systems compared to SB1, and we refer to Simón-Díaz et al. (in prep.) for further discussion.
@@ 5.3 Implications for stellar evolution models
This leaves the main difference for the position of the TAMS in Fig. 5 set by the overshooting parameter (). In the models of Brott et al. (2011), this parameter is fixed to =0.335, whereas in Ekström et al. (2012), it takes the value of =0.10. For Choi et al. (2016), they use the exponential overshoot formalism, which for models without initial rotation is equivalent to 0.20666We note that this comparison is not straightforward in models with rotation because it induces diffusive processes that contribute to chemical mixing and thus the length of the MS.. In general, in all available models, the coefficients that regulate this parameter have been set to match either the empirical location of the drop in the relative number of stars (e.g., Ekström et al. 2012), or the drop in rotational velocities (e.g., Brott et al. 2011). The latter would be more risky under the aforementioned assumption of a low initial rotation, plus the additional uncertainty in the case that fast-rotating objects have their origin in binary interaction (see Sect. 5.2). Additionally, the range of available masses used to calibrate the coefficients in those models where overshooting is not mass-dependent is an additional source of discrepancy. This explains, for example, the better agreement between the models by Brott et al. (2011) and Ekström et al. (2012) for intermediate masses below 15 M⊙.
Recent work by Martinet et al. (2021) used the observational information presented in Castro et al. (2014) to investigate the sizes of the convective cores and constrain the amount of overshooting. This and other works have also shown that the size of the cores is mass-dependent (e.g. Langer & Kudritzki 2014; Claret & Torres 2019; Tkachenko et al. 2020, 2024). Using our volume-limited sample and its high degree of completeness, we can contribute to a better definition of the width of the MS by adding new constraints to fine-tune the overshooting parameter and its dependence on the initial mass for evolutionary masses from 12 M⊙ and up to 40 M⊙. In this regard, we show in Sect. 4.3 that our empirical TAMS lies between the theoretical TAMS of Ekström et al. (2012) and Choi et al. (2016), and the one of Brott et al. (2011). Although more in-depth comparisons are certainly required, our results suggest higher overshooting values in the first two cases and a lower value above 25 M⊙ in the latter case. Moreover, the lack of an overdensity of objects with 25 – 40 M⊙ below 20 kK as shown in Castro et al. (2014) (see Sect. 4.5) also adds an important constraint for future models. Although some authors attributed the observed extension to the effect of inflation (e.g., Choi et al. 2016; Sanyal et al. 2017), this is no longer needed within that range of masses. Unfortunately, the number of stars in our sample is not sufficient to draw conclusions above 40 M⊙.
@@
The initial stellar rotation also influences the extension of the MS predicted by evolutionary models (see, e.g., Meynet & Maeder 2000; Brott et al. 2011; Martinet et al. 2021). Although, as pointed out in Sect. 4.3, we cannot directly compare our empirical FR-TAMS with the theoretical ones, our large sample compared to previous studies (e.g. Vink et al. 2010; Castro et al. 2014; McEvoy et al. 2015) allows us to withdraw one conclusion. If fast-rotating stars retain their spin rates (regardless of whether they got it as a result of binary interaction, or they were born as such), then rotationally induced effects seem to shorten the width of the MS for higher masses and extend it for lower ones. This behavior could be similar to what predicted by Ekström et al. (2012) for / = 0.4. However, in that case, the models by Ekström et al. (2012) also predict an important spin-down on the initial part of the MS, which is not consistent with our observations.

### v2

@@ 4.2 Spin-rates of the sample
Another observational characteristic proposed to delineate the location of the TAMS is the drop of the upper envelope of the distribution as a function of (see Vink et al. 2010; Brott et al. 2011). Figure 4 shows our distribution of values against separating stars in the same luminosity ranges as in Fig. 3. To qualitatively evaluate the aforementioned feature, we included all the stars regardless of their distance. The figure indicates the location of the ZAMS (gray areas), the proposed empirical TAMS (dashed colored lines), and outlines the critical rotational velocity (; solid green lines), that is, the equatorial velocity at which the centrifugal force balances gravity, as a function of (see below). We also include the median values for stars within bins of 4 kK (black-edge colored dots), and the associated 25% and 75% (solid black lines) and 10% and 90% (dashed gray lines) percentiles. Although a detailed discussion of the connection with the TAMS is presented in Sect. 5.1.1, here we highlight the three most interesting features of Fig. 4 (see also Appendix C).
A first feature is the separation of the stars into two main components (see also Fig. 10 of de Burgos et al. (2024b)): a low- densely populated group of stars (colored differently) and a less populated group of fast-rotating objects (cyan circles; indicated also in Fig. 2). Following de Burgos et al. (2024b), we adopt the value of =100 km s-1 that best separates both components. This separation is well known to be present in the O-star domain and extends toward the early-B supergiant domain (see, e.g., Holgado et al. 2022; de Burgos et al. 2023, and references therein). Using as reference those stars within 2500 pc, we find a decreasing fraction of fast-rotating objects from 50% for stars with luminosities below 4.65 dex, to 35% for stars between 5.0 and 4.65 dex, and to 30% for stars with luminosities higher than 5.0 dex. On average, 38% of the stars located on the hot side of the TAMS belong to the fast-rotating component. A further discussion of both components can be found in Sect. 5.1.2 and Sect. 5.3.
@@
A third characteristic is the consistent decrease of the maximum reached by stars with decreasing within each luminosity range. In fact, the upper envelopes closely follow , which is shown with a solid green line (shown also in the right panel of Fig. 2) and has been calculated from the same models used in Fig. 2 employing Eq. 2 in Langer (1998). All these features are discussed in more detail in Sect. 5.1.
@@ 4.4 Distribution of SB1 objects in the HR diagram
Additionally, observational biases due to the different intrinsic brightnesses of the stars are likely present in their work, whereas our volume-limited sample with high degree of completeness solves this issue. Furthermore, while the sHR diagram is used in Castro et al. (2014), here we make use of the HR diagram, combined with additional information on the spin-rate properties and binary fraction of the stars (Sects. 4.2 and 4.4). Moreover, we consider double the number of stars in Castro et al. (2014) up to 2500 pc within the same range. We also note that in Castro et al. (2014), the surface gravities were not corrected by centrifugal forces (see Herrero et al. 1992; Repolust et al. 2004), shifting fast-rotating stars toward apparently higher luminosities in the sHR diagram.
Despite the differences, Castro et al. (2014) found a drop in the density of stars below 25 M⊙ at 20 kK, which was also interpreted as the TAMS. However, for stars with 25 – 40 M⊙ the drop was located at 10 kK. As shown in Figs. 2 and 3, our results indicate that the drop occurs throughout the full range of luminosities and masses considered here, and it is located approximately at the same (24 – 22 kK). The overdensity of stars shown in Castro et al. (2014) beyond 20 kK, which is not observed in any of the panels of Fig. 2, is likely the result of a selection bias introduced by considering any star with available parameters from the literature rather than a volume-limited sample. In particular, since mid-B-type stars are intrinsically brighter in the optical compared to early-B- and O-type objects, the aforementioned overdensity is likely caused by mid-B-type stars located at larger distances.
@@ 5.1 Spin-rate properties of the sample
As indicated in Sect. 4.2, two main observable features are known to characterize the spin rates of luminous O- and B-type stars. One is the bimodal distribution of the projected rotational velocity (e.g., Conti & Ebbets 1977; Howarth et al. 1997; Dufton et al. 2013; Ramírez-Agudelo et al. 2013; Holgado et al. 2022; de Burgos et al. 2024b), which includes a main component of slowly rotating stars and a secondary more extended tail of fast-rotating objects. The latter have been proposed to be the products of binary interactions (e.g., Ramírez-Agudelo et al. 2013; de Mink et al. 2014; Ramírez-Agudelo et al. 2015; Holgado et al. 2022; Britavskiy et al. 2023; de Burgos et al. 2024b). The second feature is the lack of fast-rotating stars below a certain temperature (see Howarth et al. 1997; Ryans et al. 2002; Vink et al. 2010; Simón-Díaz & Herrero 2014; McEvoy et al. 2015; de Burgos et al. 2024b). As indicated by different studies, this characteristic could be interpreted as the empirical location of the TAMS (see Vink et al. 2010; Brott et al. 2011; Martinet et al. 2021; de Burgos et al. 2024b). We explore this possibility further.
@@ 5.1.1 The lack of fast-rotating stars beyond the TAMS
On the other hand, stars with expanding envelopes and high spin rates near the TAMS can get too close to their critical rotational velocities (see, e.g., Meynet & Maeder 2006; Meynet et al. 2007). It is predicted that before reaching that point, stars expel enough material into the medium (e.g., Kurfürst et al. 2014), triggering a spin-down of the star’s rotation (e.g., Georgy et al. 2011; de Mink et al. 2013). As shown in Fig. 4 (see also the right panel of Fig. 2), the critical rotation that we extract from the models approximates the upper envelope of the distribution, particularly in the two central panels. We note that, except for the top panel, we could, in principle, expect to find some fast-rotating objects that have not yet reached their .
@@
A third possibility is related to mass loss. In this respect, Vink et al. (2010) suggested that the bi-stability nature of the winds (see Pauldrach & Puls 1990; Vink et al. 1999, 2000) can lead to an important increase in the mass-loss rates at 22 kK, and to a subsequent enhancement of the angular momentum loss (known as the bi-stability braking). However, in de Burgos et al. (2024a) (see also Rubio-Díez et al. 2022; Krtička et al. 2024; Bernini-Peron et al. 2024) it is shown that such an increase in mass-loss rates (by a factor of ten to twenty) is not observed444Newer prescriptions, such as those of Björklund et al. (2021), do not predict any enhancement near our empirical TAMS.. Therefore, we argue that the lack of fast-rotating stars is not due to the latter proposed mechanism.
@@
Given these considerations, the lack of fast-rotating objects could be the result of the statistically and relatively lower number of stars expected beyond the TAMS, that is, the chances of observing these objects is lower. This could be further supported by the following. Considering that for stars with = 5.35 – 4.65 (where we account for more stars), the histograms of Fig. 3 show the largest difference in the relative number of stars when crossing the TAMS, which we locate at an average of 22.5 kK. For stars up to 2500 pc, there are five times fewer stars between those located on the hot and the cool sides of the TAMS within 5 kK. Considering that the proportion of fast-rotating objects located within each side with respect to the total is 1:3, this exercise would result in only six fast-rotating objects below 22.5 kK. We account for two objects that meet these conditions, which we consider sufficient to support the aforementioned possibility. We note that we do not account for the fact that some of the fast-rotating stars may not be seen as such because of a projection effect. Furthermore, if fast-rotating stars are the product of binary interaction, it is expected to occur mainly when at least one of the components of the binary system is still on the MS (see de Mink et al. 2013, 2014). Therefore, it would be difficult to observe these objects beyond the TAMS. Additionally, fast-rotating objects evolving near or beyond the TAMS will eventually reach their , thus becoming slowly rotating stars. Nonetheless, one cannot rule out the possibility of unconsidered effects or forces that can trigger moderate braking toward the end of the MS. This can include, for example, the effect of magnetic fields (see Keszthelyi et al. 2020, 2022).
@@ 5.1.2 The evolution of along the MS
Interestingly, the median values of the two top panels in Fig. 4 show similar trends toward lower effective temperatures. The values are approximately constant until 30 kK and then slowly decrease toward the cool end. This trend is less obvious in the third panel, but may still be present, and it is certainly not the case of the bottom panel, which we discuss separately in Appendix E. We also note that the first median value may not be so representative, as it overlaps with the position of the ZAMS, where the number of stars is significantly lower than for the other mean values. Excluding that one, our results strongly suggest a slow-braking process occurring along the MS, which has already been empirically supported by the works of Holgado et al. (2022); de Burgos et al. (2024b). Furthermore, the number of stars on the hot side of the TAMS with 100 km s-1 (i.e., the low- component of the spin-rate distribution) represents 65% of the total number. This large fraction may indicate that massive stars are born with low or mild initial rotations (not more than 0.2 of the critical velocity), as indicated also in the aforementioned works. However, we note that since we do not account for projection effects, that percentage may represent a lower limit.
@@ 5.2 The impact of binaries
In Sect. 4.4 we present the distribution of detected SB1 systems in an HR diagram, limited to a distance of 2500 pc. Figure 6 shows the proposed TAMS and the lack of SB1 systems beyond it, on the cool side. By evaluating the fraction of SB1 systems on each side of the TAMS, we obtain 39% within = 30 – 22.5 kK and 15% within = 22.5 – 15 kK, highlighting an important decrease when crossing the TAMS. Interestingly, the average percentage is similar to the 236% found by Dunstall et al. (2015) in the 30 Doradus region of the Large Magellanic Cloud. Furthermore, our results approach those found by McEvoy et al. (2015) in the Tarantula Nebula, where SB1 systems were not detected below a given temperature. In this work, we show that SB1 systems are also present beyond the TAMS, but their number is very low. We conclude that the lower percentage of SB1 systems beyond the TAMS can be used to infer its location with respect to . We also argue that the few SB1 systems beyond the TAMS likely correspond to stars heading to the RSg phase, since this phase practically eliminates all detectable binaries beyond the TAMS (see de Mink et al. 2014).
@@
Our results also indicate that the percentage of SB1 systems within = 31 – 22.5 kK decreases in the upper half of the HR diagram. In particular, from 40% for stars with ¡ 5.0 to 31% for stars above that luminosity. This case could be partially explained by the higher detection threshold of RV (up to 25 km s-1; see Appendix D) toward our highest luminosities, which would eliminate some true SB1 systems.
Similarly as for SB1 systems, the presence of a drop in SB2+ systems on the cool side of the TAMS could provide additional evidence of its location. Although SB2+ systems were intentionally excluded from our sample, and therefore, we do not count with reliable quantitative spectroscopic analysis of each individual component, we can still use to some extent the available information on their spectral types as a proxy of the corresponding effective temperatures.
@@
These results also provide evidence that the number of SB2+ systems in the BSG domain before reaching the TAMS is lower than for O9 stars. In this regard, given their relatively low number, the lack of SB2+ systems in the sample should not produce important biases with respect to the distribution of stars in Figs. 2, 3. However, the difference between O9- and B0-type stars could partially explain the apparent lower density of objects around 32 kK (see Sect. 4.1). We also notice the larger number of B3 – B4 SB2+ systems compared to SB1 and we refer to Simón-Díaz et al. (in prep.) for further discussion.
@@ 5.3 Implications for stellar evolution models
This leaves the main difference for the position of the TAMS in Fig. 5 set by the overshooting parameter (). In the models of Brott et al. (2011), this parameter is fixed to =0.335, whereas in Ekström et al. (2012), it takes the value of =0.10. For Choi et al. (2016), they used the exponential overshoot formalism, which for models without initial rotation is equivalent to 0.20666We note that this comparison is not straightforward in models with rotation because it induces diffusive processes that contribute to chemical mixing and thus the length of the MS.. In general, in all available models, the coefficients that regulate this parameter have been set to match either the empirical location of the drop in the relative number of stars (e.g., Ekström et al. 2012), or the drop in rotational velocities (e.g., Brott et al. 2011). The latter would be more risky under the aforementioned assumption of a low initial rotation, plus the additional uncertainty in the case that fast-rotating objects have their origin in binary interaction (see Sect. 5.2). Additionally, the range of available masses used to calibrate the coefficients in those models where overshooting is not mass-dependent is an additional source of discrepancy. This explains, for example, the better agreement between the models by Brott et al. (2011) and Ekström et al. (2012) for intermediate masses below 15 M⊙.
Recent work by Martinet et al. (2021) used the observational information presented in Castro et al. (2014) to investigate the sizes of the convective cores and constrain the amount of overshooting. This and other works have also shown that the size of the cores is mass-dependent (e.g., Langer & Kudritzki 2014; Claret & Torres 2019; Tkachenko et al. 2020, 2024). Using our volume-limited sample and its high degree of completeness, we can contribute to a better definition of the width of the MS by adding new constraints to fine-tune the overshooting parameter and its dependence on the initial mass for evolutionary masses from 12 M⊙ and up to 40 M⊙. In this regard, we show in Sect. 4.3 that our empirical TAMS lies between the theoretical TAMS of Ekström et al. (2012) and Choi et al. (2016), and the one of Brott et al. (2011). Although more in-depth comparisons are certainly required, our results suggest higher overshooting values in the first two cases and a lower value above 25 M⊙ in the latter case. Moreover, the lack of an overdensity of objects with 25 – 40 M⊙ below 20 kK as shown in Castro et al. (2014) (see Sect. 4.5) also adds an important constraint for future models. Although some authors attributed the observed extension to the effect of inflation (e.g., Choi et al. 2016; Sanyal et al. 2017), this is no longer needed within that range of masses. Unfortunately, the number of stars in our sample is not sufficient to draw conclusions above 40 M⊙.
@@
The initial stellar rotation also influences the extension of the MS predicted by evolutionary models (see, e.g., Meynet & Maeder 2000; Brott et al. 2011; Martinet et al. 2021). Although, as pointed out in Sect. 4.3, we cannot directly compare our empirical FR-TAMS with the theoretical ones, our large sample compared to previous studies (e.g., Vink et al. 2010; Castro et al. 2014; McEvoy et al. 2015) allows us to withdraw one conclusion. If fast-rotating stars retain their spin rates (regardless of whether they got it as a result of binary interaction, or they were born as such), then rotationally induced effects seem to shorten the width of the MS for higher masses and extend it for lower ones. This behavior could be similar to what predicted by Ekström et al. (2012) for / = 0.4. However, in that case, the models by Ekström et al. (2012) also predict an important spin-down on the initial part of the MS, which is not consistent with our observations.